Please provide an explanation for the following

A. During 2018, Fossil INC had sales of $600,000, cost of goods sold of $450,000, administrative and selling expenses of $95,000, depreciation expense of $140,000 and interest expense of $70,000. The tax rate is 35 percent. Ignore any tax loss carry back or carry forward provisions. What is the operating cash flow for Fossil?

B. If Fossil, Inc. has sales of 12,300, total assets of 13,000, a debt to equity ratio of 1.3 and a return on equity of 15 percent, what is Fossils net income?
Answer #1

A: Sales = 600000

Less: Cost of goods sold = 450000

Selling and administrative expenses = 95000

Depreciation 140000

EBIT = -85000

Less tax = 35% = -29750

Operating income after tax = -55250

Add depreciation 140000

Operating cash flow = 84750

2: Debt to equity = 1.3/1

Hence debt = 1.3/2.3 of total assets

Equity = 1/2.3 of total assets

Equity = 13000*1/2.3=5652.174

ROE = 15% = Net Income/ Equity

Net Income = 15%*5652.174

= 847.83